OpenStudy (anonymous):

PLEASE HELP ASAP! Which of the following is the solution to the equation sqrt 5 ^8n = 125 ^ (n + 3) ?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

9

jhonyy9 (jhonyy9):

sqrt5 is 5 on exponent1/2 than on left part will be 5 on exponent 8n/2 = 4n so than on the right part will be 5 on exponent 3(n+3) - than 4n=3(n+3) 4n=3n+9 n=9

jhonyy9 (jhonyy9):

sorry for this firstly answer but than i have missed the sqrt of 5